OpenStudy (anonymous):

Four times the sum of a number and 15 is at least 120. Let x represent the number. Find all possible values for x. A. x ³ 26 B. x ³ –15 C. x ³ 15 D. x ³ –26

OpenStudy (anonymous):

i know it is not (A)

OpenStudy (anonymous):

i think it is A

OpenStudy (anonymous):

4(x+15)>=120

hartnn (hartnn):

what does 3 mean there?

OpenStudy (anonymous):

^3 X*X*X*=

hartnn (hartnn):

then u need to verify your options because 4(x+15)>=120 gives x+15>=30 or x>=15 is there any such option??

OpenStudy (anonymous):

(C)

hartnn (hartnn):

ok then.

OpenStudy (anonymous):

thankx =)

hartnn (hartnn):

welcome :)